OpenStudy (anonymous):

What is the product? 5X(- 6) x ( - 8) x (2) x ( -1) A. –480 B. –240 C. 240 D. 480

OpenStudy (anonymous):

@YoungStudier

OpenStudy (mathstudent55):

Count the number of negative signs. If it is even, the answer is positive. If it is odd, the answer is negative. Now ignore the signs, multiply the numbers with no signs, and just write the correct sign according to the above rule.

OpenStudy (mathstudent55):

\( 5 \times (\color{red}{-} 6) \times ( \color{red}{-} 8) \times (2) \times ( \color{red}{-}1)\) How many negative signs are there?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

You're welcome.